Minutes of the Trigger Operation Meeting (, 2007)
Attendees: Debbie Bard, Fang Fang, Rahmat Rahmat.
- The detector is starting to run as per-normal again. This week has seen a couple of pages for minor things, like L3 reference plots being off. Paged Friday night for damage of ~1000 events (see LiveFastMon plots) but was paged after run had finished and problem had cleared with the start of the new run. ODF log file showed out-of-order and out-of-synch errors but no spurious interrupt messages.
- See email from Mark here.
- Fang has been working on the teststand and has connected the ZPD to the GLT, as the ZPD power supply is now fixed. One cannot, therefore, use the teststand GLT in the current state - there are instructions attatched to the board on how to reset it if it's needed in the coming days.
- The problem with the memory tests has been isolated in the code and Fang has a fix which works at the teststand.
Data Quality and Monitoring
- DQ checks of Run 6 and reprocessing continues and so far so good.
- The problem with the stripcharts has been isolated - in Italy the charts cannot be made with lines between the points, as required by the trigger plots. The connecting lines have therefore been removed in the Italian code, and Rahmat must re-mmake the trigger stripcharts himself every time.
This page is created by Debbie Bard (firstname.lastname@example.org); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.